Every year the Earth travels about 1x10*9 km as it orbits the Sun. What is Earth’s average speed in km/h?
1
Expert's answer
2020-09-16T10:11:31-0400

The average speed of Earth

"v=\frac{2\pi R}{T}=\frac{2\pi\times 10^9\: \rm km}{365\times 24\:\rm h}=7.2\times 10^5\:\rm km/h"